Draw the structural formula for these compounds.

3,4-dimethylhexyne

2,2-dimethyl-5-ethyl-3-heptene

ethylcyclopentene
1
Expert's answer
2012-10-17T11:16:23-0400
HC=C-HC(CH3)-HC(CH3)-CH2-CH3

H3C-C(CH3)2-CH=CH-CH(C2H5)-CH2-CH3

CH3-CH2-(C5H7) (C5H7 is 5 carbons ring with one double bond between two of carbons)